# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ 5.4 Examples
$\frac{{\cal Z}_{\rm CT}[V_{4}](C_{2}\otimes{\mathbb{1}}^{\otimes 2})}{{\cal Z}_{\rm CT}[0](C_{2}\otimes{\mathbb{1}}^{\otimes 2})}=(1-\lambda)^{-\frac{1}{2}(N^{2}-1)^{2}}(1-(1+N)\lambda)^{-(N^{2}-1)}(1-(1+2N)\lambda)^{-\frac{1}{2}}=\frac{{\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},Y[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]}{{\cal Z}_{\rm\hat{H}T}[0,Y[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]}\;.$ (5.65)
@@
Figure 11: The couplings $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ correspond to the same bubble with different assignments of colors. The same applies to the couplings $\lambda_{4}$, $\lambda_{5}$, and $\lambda_{6}$. The coupling $\lambda_{7}$ corresponds to a vertex with the topology of a complete graph. In all cases, we highlight the strand of color 1 in red.
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{6},\hat{Y}[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes 2}})}{\cal D}\hat{\Phi}\;e^{-\frac{N}{2}{\rm Tr}(\hat{\Psi}^{2})+\hat{V}_{6}(\hat{\Psi})}\;$, (5.68)
@@
Figure 12: In the effective self-adjoint model, most vertices have different topologies, except for the pair $(\lambda_{2},\lambda_{3})$ and the pair $(\lambda_{5},\lambda_{6})$.
Equivalences for trace-invariant potentials.
[…]
Specific examples of these trace-invariant potentials can be found in
  • (i)
  • (ii)
See also the corresponding Figures for (i) Figures 9 and 10, and (ii) Figures 11 and 12 for the graphic illustrations of these trace-invariant potentials.
@@
Figure 13: Examples of the interaction bubbles present in left: the complex tensor model with potential (5.75) and right: the self-transpose tensor model with potential (5.76) for the case $D=3$. In this case, the thinner red lines correspond to the indices contracted with $\varphi$ and $\varphi^{\dagger}$.
@@
An example of the type of vertices in this model is shown in Figure 13. Next, we consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function with potential (5.75)
@@
${\cal Z}_{\rm HT}[{V}_{\rm 4},Y[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}\left(\Psi\Phi\right)-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)+V_{4}(\Phi)}\;,$ (5.78)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm HT}[V_{4},Y[R]]}{{\cal Z}_{\rm HT}[0,Y[R]]}\;.$ (5.79)
In this example the enhanced quartic pillow interaction on the order–$D$ complex tensor is mapped to a quadratic interaction on the order–$2D$ self–adjoint tensor, providing a particularly simple illustration of how bubble structure changes under the equivalence.
@@ 5.5 A prospect on real tensor models
Up to this point, we have shown equivalences of complex tensor models only. However, real tensor models are of considerable interest for quantum gravity, and several families have been extensively studied in the physics literature [Eichhorn:2018phj, Eichhorn:2018ylk, Eichhorn:2019hsa]. One motivation is that the numerical values of the critical exponents characterizing the continuum limit appear to be insensitive to whether the tensor ensemble is complex or real. Since the real case is technically simpler, such models are often preferred in FRG. Nevertheless, although some results have suggested the existence of a continuum limit compatible with four-dimensional quantum gravity [Eichhorn:2018phj, Eichhorn:2018ylk, Eichhorn:2019hsa], no study has yet demonstrated the emergence of geometries beyond the branched–polymer phase. This limitation has renewed interest in introducing a causal structure, in close analogy with what has been achieved in the causal matrix model case [Benedetti:2008hc, Castro:2020dzt]. It is precisely in this context that the equivalences established in this work may provide a useful framework for implementing such constraints.
@@
This result shows that a tensor model implementing a causal-like [^11] However, this model generates graphs that are strictly a subset of CDT as their dual triangulation just like the model presented in Subsection 4.1. constraint can be reformulated as a model in which this constraint is traded for a symmetry of the tensor indices (namely, symmetry under transposition). As a consequence, this equivalence offers a setting in which the continuum limit may be investigated using FRG techniques without having to deal explicitly with the presence of the constant matrix $C_{2}$. Whether this reformulation is technically simpler in practice remains an open question and is left for future investigation. Moreover, in the tensor models considered here, the interaction are of tetrahedral type. This structure is more closely aligned with the geometric picture of gluing $D$-simplices to construct a discrete spacetime, thereby their connection with simplicial approaches to quantum gravity is more straightforward.
Given this motivation, we move on to probing the equivalence by computing the partition functions of both models. We do this for the case $D=3$, however, we expect that this equivalence should be generalized easily to any $D$.
@@
are the remaining eigenvectors of $\Sigma$ with eigenvalues $\lambda$. These are also eigenvectors of $I_{N}\otimes I_{N}+\lambda\Sigma$, $u_{ij}$, with eigenvalues $1+\lambda$, and $v_{ij}$ with eigenvalues $1-\lambda$. Therefore, we have that
@@
Figure 14: An example of a segment of a graph in the order-$3$ real tensor model. The red lines indicate the presence of the matrix $C_{2}$ in the propagator. In this case, the interactions are tetrahedral. In general, any graph in this model will be a closure of a sequence of vertices as presented here.
Figure 15: An example of a segment of a graph in the effective model: a real self-transpose tensor model without $C_{2}$. The red boxes correspond to the interactions in this model. The parallel black lines represent the propagators. Any graph in this model will be a closure of a sequence of vertices as presented here.
In summary, the proved equivalence shows that a real tensor model of order 3 with a rigidity constraint is equivalent to a real self-transpose tensor model of order 4 without rigidity constraints. In terms of the Feynman graph expansion, the rigidity constraint in the order-$3$ model leads to necklace/chain-type graphs as illustrated in Figure 14. This is reminiscent of the complex matrix model (3.9). In other words, the tetrahedral interactions deform into necklace/chain-type graphs.
@@ 6 Conclusions and outlook
In this paper, we have explored the equivalences between complex and self-adjoint random matrix and tensor models. More specifically, we established exact equivalences via intermediate–field representations between
@@
  • complex matrix models with a general covariance $(P,Q)$ and potentials depending on $M^{\dagger}M$, and self–adjoint two–matrix models whose intermediate field carries a logarithmic dually-weighted potential;
@@
  • complex tensor models with covariance $R$ and potentials depending on $\varphi\otimes\varphi^{\dagger}$, and self–adjoint tensor models with logarithmic potentials.
In all cases the equivalence holds at the level of partition functions and all trace invariants built from the self–adjoint combination (either $M^{\dagger}M$ or $\varphi\otimes\varphi^{\dagger}$). We summarize more concretely our main results below.
Matrix models.
$\mathcal{Z}_{\rm HM}[V,Y]=\iint_{{\rm H}(N)^{2}}{\cal D}A\,{\cal D}B\;e^{-iN{\rm Tr}(AB)+Y[P,Q](B)+V(A)},\quad Y[P,Q](B)=-{\rm Tr}\ln\left(\mathbb{1}^{\otimes 2}-i\,Q\otimes(PB)\right).$ (6.2)
Then, Theorem 4.3 implies that: for any potential $V$,
$\frac{\mathcal{Z}_{\rm CM}[V](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V,Y[P,Q]]}{\mathcal{Z}_{\rm HM}[0,Y[P,Q]]}.$ (6.3)
Our central result is that this agreement not only holds at the level of Feynman expansions, but also the two partition functions are exactly equivalent via an intermediate–field representation.
Equivalently, any observable depending only on the self–adjoint combination $M^{\dagger}M$ can be computed in either model. This is stated in the following corollary. Equivalently, Corollary 4.3.1 states that, for any potential $V$ and any function $f:{\rm H}(N)\to\mathbb{C}$,
$\left\langle f(M^{\dagger}M)\right\rangle[V](P,Q)=\left\langle f(A)\right\rangle[V,Y[P,Q]].$ (6.4)
Several specializations follow from this equivalence given in Theorem 4.3 and Corollary 4.3.1.
  • For dually weighted choices of $V$, the complex matrix model $\mathcal{Z}_{\rm CM}$ is equivalent to a self–adjoint two–matrix model $\mathcal{Z}_{\rm HM}$ in which both matrices carry dually weighted potentials (Section 4.4).
  • For $Q=C_{2}$ and a suitable quadratic potential $V$, the complex matrix model $\mathcal{Z}_{\rm CM}$ becomes equivalent to a purely Gaussian self–adjoint one–matrix model (cf. (4.40)).
  • For the generalized choice $Q=C_{k}$, the equivalence enforces a decomposition of one family of index cycles into disjoint $k$-blocks, inducing a natural partition–matroid structure (4.44).
Tensor models.
The equivalence between complex and self-adjoint theories persists also in tensor models. Let $\varphi\in{\mathbb{C}}_{N}^{\otimes D}$ be a random complex order–$D$ tensor with covariance $R\in{\mathbb{C}}_{N}^{\otimes D}\otimes{\mathbb{C}}_{N}^{*\otimes D}$, and $V(\varphi\varphi^{\dagger})$, a scalar-$U(1)$ invariant potential, as elaborated in Subsection 5.1. The partition function of the complex theory in question is (5.8)
We consider random tensors $\Phi,\Psi\in{\rm H}({\mathbb{C}}_{N}^{\otimes D})$ distributed according to the partition function of the self-adjoint theory (5.15) with potential (5.24)
${\cal Z}_{\rm HT}[V,Y[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}(\Phi\Psi)+Y[R](\Psi)+V(\Phi)}\,,\qquad Y[R](\Psi)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)\,$, (6.6)
where the potential $Y[R](\Psi)$ gives $\Psi$ the role of the dually weighted intermediate field. Our Theorem 5.3 states that, for any potential function $V$,
$\frac{\mathcal{Z}_{\rm CT}[V](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm HT}[V,Y[R]]}{\mathcal{Z}_{\rm HT}[0,Y[R]]}\;.$ (6.7)
Equivalently, as stated in Corollary 5.3.1, for any potential $V$ and any function $f:{\rm H}({\mathbb{C}}_{N}^{\otimes D})\to\mathbb{C}$,
$\left\langle f(\varphi\varphi^{\dagger})\right\rangle[V](R)=\left\langle f(\Phi)\right\rangle[V,Y[R]].$ (6.8)
Under an additional symmetry, i.e., the partial-$U(N)$ invariance explained in Subsection 5.1, the self–adjoint tensor model can be reduced to a model with tensors of order 2(D1)2(D-1) (5.38). By working with the partial trace ${\rm Tr}_{(1)}(\varphi\varphi^{\dagger})$ under its equivalence with a certain order-$D$ partially traced complex tensor model, (5.37) as achieved in Theorem 5.4. This takes the form
$\frac{\mathcal{Z}_{\rm CT}[\hat{V}\circ{\rm Tr}_{(1)}](R)}{\mathcal{Z}_{\rm CT}[0](R)}=\frac{\mathcal{Z}_{\rm\hat{H}T}[\hat{V},\hat{Y}[R]]}{\mathcal{Z}_{\rm\hat{H}T}[0,\hat{Y}[R]]}\;.$ (6.9)
In Section 5.4, we gave some examples of the equivalences achieved in Theorems 5.3 and 5.4, which we revisit below.
For Theorem 5.3,
  • The quartic enhanced pillow tensor model of order $D$ (5.75) is mapped to an order–$2D$ self–adjoint tensor model with a purely quadratic effective potential (5.76).
  • The quartic pillow tensor model of order $D$ (5.54) is mapped to an order–$2D$ self–adjoint tensor model with a purely quadratic effective potential (5.55).
For Theorem 5.4,
  • The quartic pillow tensor model of order $D$ (5.59) is mapped to an order–2(D1)2(D-1) self–adjoint tensor model with a purely quadratic effective potential (5.60).
  • An order–33 complex sextic tensor model (5.66) with potential (5.67) is mapped to an order-$4$ self–adjoint cubic tensor model (5.68) with potential (5.69).
  • For factorized propagators in one of the colors i.e. $R=Q\otimes\hat{P}$, we get a weighting of vertices (5.49), similarly to the dually weighted matrix models in Subsection 3.1. If $Q=C_{2}$, the tensor equivalence reproduces a self-adjoint one-tensor model, in direct parallel with the matrix model examples.
Notably, the first three are examples where a non–Gaussian tensor model is shown to be equivalent to a Gaussian one.
We remark that when $D=2$ and the propagator factorizes as $R=Q\otimes P^{{\rm T}}$, the tensor equivalence of Theorem 5.4 specializes to the matrix equivalence of Theorem 4.3 for interactions with partial-$U(N)$ invariance. In this setting, the complex matrix model is naturally rewritten as an order–$2$ tensor model, as in (5.48), making the matrix equivalence a direct corollary of the tensor result.
In summary, our results show that complex matrix ($M$) and tensor ($\varphi$) models with nontrivial covariance and interactions depending on $M^{\dagger}M$ or $\varphi\otimes\varphi^{\dagger}$ admit exact self–adjoint representations in which the covariance data is carried by a logarithmic intermediate field. This equivalence provides both a conceptual reinterpretation of such models and a practical computational tool: any expectation values/observable built from the self–adjoint combination may be computed in whichever formulation (complex or self–adjoint).
Relation to existing intermediate field representations.
where ${\cal N}=(\mathcal{Z}_{\rm HT}[0,Y[R]])^{-1}(\mathcal{Z}_{\rm CT}[0](R))$, and with additional partial symmetry (Theorem 5.4),
for when there is a $\hat{V}$ such that $V(\varphi\varphi^{\dagger})=\hat{V}({\rm Tr}_{(1)}(\varphi\varphi^{\dagger}))$. The above equalities are exact, for any covariance $R$ and any potential $V$. Previously known intermediate field representations arise from suitable specializations of this identity (6.10).
    In [Bonzom2015ColoredTO], the intermediate field formula is developed for a single trace-invariant tensor interaction ${\rm Tr}_{B}(T,\bar{T})$ with unit covariance and is proven as a formal series in the coupling constant $\lambda$. In our notation, their intermediate field formula corresponds to choosing $R={\mathbb{1}}$ and $V(\Phi)=\frac{\lambda}{n}N^{s}\,V_{B,\Omega}(\Phi)$ (Eq. (52) in [Bonzom2015ColoredTO])with $\varphi$ being the flattening of the tensor $T$. Making these substitutions in (6.10), we obtain the following two-field representation (Eq. (53) in [Bonzom2015ColoredTO])
    $Z_{B}(\lambda,N)=\iint{\cal D}\Phi\,{\cal D}\Psi\;e^{-i{\rm Tr}(\Psi\Phi)-{\rm Tr}\ln({\mathbb{1}}^{\otimes D}-i\Psi)+\frac{\lambda}{n}N^{s}V_{B,\Omega}(\Phi)}.$ (6.12)
    Formally integrating out the intermediate field $\Psi$ yields precisely the formulation that appears in [Bonzom2015ColoredTO]. However, a conceptual difference is that our equality holds exactly, while the derivation in [Bonzom2015ColoredTO] is performed at the level of formal power series.
  • Hubbard–Stratonovich representation.
    The classical intermediate field representation for the quartic one-matrix model, see e.g. [gurau2014analyticityresultscumulantsrandom], follows from choosing $D=2$, $R=\sqrt{\frac{\lambda}{N}}{\mathbb{1}^{\otimes 2}}$, and $\hat{V}(\hat{\Phi})=-\frac{1}{2}{\rm Tr}(\hat{\Phi}^{2})$. Our identity (6.11), after integrating out $\Phi$, yields
In summary, the various intermediate field constructions developed in the matrix and tensor literature arise as specific choices of the covariance $R$ and the potential $V$ in the universal identity above (6.10) that we discovered in this paper. This provides a unified analytic framework that extends these previous results to completely general potentials and covariances, without relying on perturbative expansions.
@@
From a mathematical point of view, it would be interesting to understand whether the rigid block structures arising in the $Q=C_{k}$ models admit a more intrinsic combinatorial characterization, and whether the associated random matrix and tensor ensembles can be systematically related to well–defined classes of matroidal or hypergraph structures [oxley2006matroid].
@@
A point that remains to be explored concerns the convergence conditions for the tensor $R$, which plays the role of the covariance in the tensor model case. This analysis should be analogous to the one presented in Appendix B for the covariance $(P,Q)$ in the matrix model case. We plan to investigate this question in future work.
@@
On the physics side, there remains to be investigated if the matrix and tensor models including the matrix $C_{2}$ we introduced, belong to universality classes interesting for quantum gravity. As we discussed, the fact that the graph structure of the models we introduced reduces to chain-like graphs does not discard them. As it has been shown [Eichhorn:2018phj, Eichhorn:2018ylk, Eichhorn:2019hsa], even when the continuum limit of some tensor models is dominated by melonic graphs, which in principle do not resemble extended geometries, their critical exponents are compatible with those of some QG models (e.g. unimodular gravity). This may be analogous to what happens in two-dimensional CDT [Benedetti:2008hc, Castro:2020dzt] where one of the critical exponents of the matrix model (namely, the string susceptibility) corresponds to that of the tree universality class. However, this tree structure does not represent the full two-dimensional geometry per se; rather, it encodes the global foliation structure. We plan to explore this research direction by carrying out an FRG analysis of the real tensor model (5.80).
@@
In this work, we decided to focus on the case $Q=C_{k}$ and, in particular, $Q=C_{2}$ as defined in (3.6). However, in [Benedetti:2008hc] $C_{2}$ was found to be suitable to impose causality in the graph expansion of Hermitian matrix models in particular. Therefore, it remains to be investigated if suitable tensor structures analogous to $C_{2}$ that serve to impose a causal structure in complex matrix and tensor models exist. If they do, it will be of interest to explore whether the equivalences proved in this work can be applied to them. Indeed, the equivalences we construct are sufficiently general to accommodate a broad class of covariance structures $(P,Q)$ in the matrix case and $R$ in the tensor model case. This is an avenue we hope to explore in future work.
@@
On the more speculative side, we noted that in the matrix model case, the dual graphs Figure 2 have a strong resemblance with layered posets (KR-orders) appearing in the Causal Set configuration space [Carlip:2022nsv]. These correspond to spacetimes with only three time steps (layers), which dominate the Causal Set path integral. They are therefore regarded as problematic, since such spacetimes do not resemble our extended universe. It remains an open question whether the continuum limit of the random matrix model (3.10) presented here retains these pathological configurations. We expect this analysis can be done using FRG techniques similarly to [Eichhorn:2013isa, Castro:2020dzt].
Acknowledgments

### v2

@@ 6.4 Examples
$\frac{{\cal Z}_{\rm CT}[V_{4}](C_{2}\otimes{\mathbb{1}}^{\otimes 2})}{{\cal Z}_{\rm CT}[0](C_{2}\otimes{\mathbb{1}}^{\otimes 2})}=(1-\lambda)^{-\frac{1}{2}(N^{2}-1)^{2}}(1-(1+N)\lambda)^{-(N^{2}-1)}(1-(1+2N)\lambda)^{-\frac{1}{2}}=\frac{{\cal Z}_{\rm\hat{H}T}[\hat{V}_{4},\hat{Y}_{\ln}[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]}{{\cal Z}_{\rm\hat{H}T}[0,\hat{Y}_{\ln}[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]}\;.$ (6.72)
@@
Figure 19: Stranded representation of the sextic vertices generated by the partition function (6.73). The couplings $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ correspond to the same bubble with different assignments of colors. The same applies to the couplings $\lambda_{4}$, $\lambda_{5}$, and $\lambda_{6}$. The coupling $\lambda_{7}$ corresponds to a vertex with the topology of a complete graph.
@@
${\cal Z}_{\rm\hat{H}T}[\hat{V}_{6},\hat{Y}_{\ln}[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]=\iint_{{\rm H}({{\mathbb{C}}_{N}^{\otimes 2}})}{\cal D}\hat{\Phi}\;e^{-\frac{N}{2}{\rm Tr}(\hat{\Psi}^{2})+\hat{V}_{6}(\hat{\Psi})}\;$, (6.75)
@@
Figure 20: Stranded representation of the cubic vertices of the intermediate effective field $\Phi$ generated by the partition function (6.75).
It is worth noticing that there is an intermediate action where all fields are present. In this sextic model, it is equal to
${\cal Z}_{\rm INT}[\hat{V}_{6},\hat{Y}_{\ln}[C_{2}\otimes{\mathbb{1}}^{\otimes 2}]]=\iint{\cal D}\varphi^{\dagger}{\cal D}\varphi\iint{\cal D}\hat{\Phi}{\cal D}\hat{\Psi}\;e^{-\varphi^{\dagger}\left(R^{-1}-i\,{\mathbb{1}}\otimes\hat{\Psi}\right)\varphi-i{\rm Tr}\left(\hat{\Psi}\hat{\Phi}\right)+\hat{V}_{6}(\hat{\Phi})}\;$, (6.77)
and incorporates the transformations shown in Fig. 22.
Figure 21: The interaction vertex $\varphi^{\dagger}\Psi\varphi$, which appears in (6.77).
Figure 22: Examples of the intermediate steps (according to (6.77)) in the equivalence between the sextic vertex represented in Fig. 19(a), Fig. 19(f), Fig. 19(g), and the cubic vertex represented in Fig. 20(a), Fig. 20(f), Fig. 20(g) ((a), (f) and (g), respectively). Intermediate fields $\Psi$ and $\Phi$ are presented in gray lines as in Figure 14.
@@
Figure 23: Examples of the interaction bubbles present in left: the complex tensor model with quartic potential (6.78) and right: the self-transpose tensor model with quadratic potential (6.79) for the case $D=3$. In this case, the thinner red lines correspond to the indices contracted with $\psi$ and $\psi^{\dagger}$.
@@
An example of the type of vertices in this model is shown in Figure 23. We consider a random complex tensor $\varphi\in\mathbb{C}_{N}^{\otimes D}$ with partition function with potential (6.78)
@@
${\cal Z}_{\rm HT}[{V}_{\rm 4},{Y}_{\ln}[R]]=\iint{\cal D}\Phi{\cal D}\Psi\;e^{-i{\rm Tr}\left(\Psi\Phi\right)-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes D}-iR\Psi\right)+V_{4}(\Phi)}\;,$ (6.81)
@@
$\frac{{\cal Z}_{\rm CT}[V_{4}](R)}{{\cal Z}_{\rm CT}[0](R)}=\frac{{\cal Z}_{\rm HT}[V_{4},{Y}_{\ln}[R]]}{{\cal Z}_{\rm HT}[0,{Y}_{\ln}[R]]}\;.$ (6.82)
In this example the enhanced quartic pillow interaction on the order–$D$ complex tensor is mapped to a quadratic interaction on the order–$2D$ self–adjoint tensor.
@@ 6.5 A prospect on real tensor models
Up to this point, we have shown equivalences of complex tensor models only. However, real tensor models are of considerable interest for quantum gravity, and several families have been extensively studied in the physics literature [18, 17, 20]. One motivation is that the numerical values of the critical exponents characterizing the continuum limit appear to be insensitive to whether the tensor ensemble is complex or real. Since the real case is technically simpler, such models are often preferred in FRG. Nevertheless, although some results have suggested the existence of a continuum limit compatible with four-dimensional quantum gravity [18, 17, 20], no study has yet demonstrated the emergence of geometries beyond the branched–polymer phase. This limitation has renewed interest in introducing a causal structure, in close analogy with what has been achieved in the causal matrix model case [7, 12]. It is precisely in this context that the equivalences established in this work may provide a useful framework for implementing such constraints.
@@
This result shows that a tensor model implementing a causal-like [^13] However, this model generates graphs that are strictly a subset of CDT as their dual triangulation just like the model presented in Subection 5.1. constraint can be reformulated as a model in which this constraint is traded for a symmetry of the tensor indices (namely, symmetry under transposition). As a consequence, this equivalence offers a setting in which the continuum limit may be investigated using FRG techniques without having to deal explicitly with the presence of the constant matrix $C_{2}$. Whether this reformulation is technically simpler in practice remains an open question and is left for future investigation. Moreover, in the tensor models considered here, the interaction are of tetrahedral type. This structure is more closely aligned with the geometric picture of gluing $D$-simplices to construct a discrete spacetime, thereby their connection with simplicial approaches to quantum gravity is more straightforward.
Given this motivation, we move on to probing the equivance by computing the partition functions of both models. We do this for the case $D=3$, however, we expect that this equivalence should be generalized easily to any $D$.
@@
are the remaining eigenvectors of $\Sigma$ with eigenvalues $\lambda$. These are also eigenvectors of $I_{N}\otimes I_{N}+\lambda\Sigma$, $u_{ij}$, with eigenvalues $1+\lambda$, and $v_{ij}$ with eigenvalues $1-\lambda$. Thefore, we have that
@@
Figure 24: An example of a segment of a graph in the order-$3$ real tensor model. The red lines indicate the presence of the matrix $C_{2}$ in the propagator. In this case, the interactions are tetrahedral. In general, any graph generated by the partition function (6.83) will be a closure of a sequence of vertices as presented here.
Figure 25: An example of a segment of a graph in the effective model: a real self-transpose tensor model without $C_{2}$. The red boxes correpond to the quadratic interactions in (6.94). The parallel dashed gray lines represent the propagators. Any graph in this model will be a closure of a sequence of vertices as presented here.
In summary, the proved equivalence shows that a real tensor model of order 3 with a rigidity constraint is equivalent to a real self-transpose tensor model of order 4 without rigidity constraints. In terms of the Feynman graph expansion, the rigidity constraint in the order-$3$ model leads to necklace/chain-type graphs as illustated in Figure 24. This is remiscent of the complex matrix model (3.9). In other words, the tetrahedral interactions deform into necklace/chain-type graphs.
@@ 7 Conclusions and outlook
In this paper, we have explored and presented newly found equivalences between complex and self-adjoint random matrix and tensor models. More specifically, we established exact equivalences via intermediate–field representations between
@@
  • complex matrix models with a general covariance $(P,Q)$ and potentials depending on $M^{\dagger}M$, and self–adjoint two–matrix models whose intermediate field carries a logarithmic dually-weighted potential (Theorem 5.3);
@@
  • complex tensor models with covariance $R$ and potentials depending on $\varphi\otimes\varphi^{\dagger}$, and self–adjoint tensor models with dually-weighted logarithmic potentials (Theorems 6.3, and 6.4).
In all cases, the equivalence holds at the level of partition functions and all trace invariants built from the self–adjoint combinations: $M^{\dagger}M$ for matrices, and $\varphi\otimes\varphi^{\dagger}$ for tensors. In general, intermediate field representation reduces the order of the polynomial interaction potential and can provide a simpler analysis. As an application of the matrix equivalence, we showed in Section 5.4 that a class of non-Gaussian matrix models can be reformulated as Gaussian ones. By focusing on the equivalence of observables, including disc functions, we obtained a new representation of the Catalan numbers.
In the random matrix model literature, a widely used technique is the character expansion, which expresses partition functions and observables in terms of irreducible characters of the symmetry group, often reducing the effective degrees of freedom and simplifying the analysis. In the present setting, Section 5.5 shows that the expectation values obtained in this work provide the necessary input to evaluate $\langle\chi_{\lambda}(A)\rangle$ for arbitrary partitions $\lambda$, thereby connecting our results to standard character-based methods.
@@
From a mathematical point of view, it would be interesting to understand whether the rigid block structures arising in the $Q=C_{k}$ models admit a more intrinsic combinatorial characterization, and whether the associated random matrix and tensor ensembles can be systematically related to well–defined classes of matroidal or hypergraph structures [32].
@@
A point that remains to be explored concerns the convergence conditions for the tensor $R$, which plays the role of the covariance in the tensor model case. This analysis should be analogous to the one presented in Appendix B for the covariance $(P,Q)$ in the matrix model case. We plan to investigate this question in future work. Furthermore, intermediate field representation played a crucial role [29] to prove Borel summability (which then uniquely defines the theory non-perturbatively) of the vector, matrix, and tensor models of quartic interaction. Therefore, our novel intermediate field representation techniques which now cover more general (i.e., dually-weighted) models may allow to proe Borel summability of a larger class of models.
@@
On the physics side, there remains to be investigated if the matrix and tensor models including the matrix $C_{2}$ we introduced, belong to universality classes interesting for quantum gravity. As we discussed, the fact that the graph struture of the models we introduced reduces to chain-like graphs does not discard them. As it has been shown [18, 17, 20], even when the continuum limit of some tensor models is dominated by melonic graphs, which in principle do not resemble extended geometries, their critical exponents are compatible with those of some QG models (e.g. unimodular gravity). This may be analogous to what happens in two-dimensional CDT [7, 12] where one of the critical exponents of the matrix model (namely, the string susceptibility) corresponds to that of the tree universality class. However, this tree structure does not represent the full two-dimensional geometry per se; rather, it encodes the global foliation structure. We plan to explore this research direction by carrying out an FRG analysis of the real tensor model (6.83).
@@
In this work, we decided to focus on the case $Q=C_{k}$ and, in particular, $Q=C_{2}$ as defined in (3.6). However, in [7] $C_{2}$ was found to be suitable to impose causality in the graph expansion of Hermitian matrix models in particular. Therefore, it remains to be investigated if suitable tensor structures analogous to $C_{2}$ that serve to impose a causal structure in complex matrix and tensor models exist. If they do, it will be of interest to explore whether the equivalences proved in this work can be applied to them. Indeed, the equivalences we construct are sufficiently general to accommodate a broad class of covariance structures $(P,Q)$ in the matrix case and $R$ in the tensor model case. This is an avenue we hope to explore in future work.
@@
On the more speculative side, we noted that in the matrix model case, the dual graphs Figure 2 have a strong resemblence with layered posets (two-orders) appearing in the Causal Set configuration space [11]. These correspond to spacetimes with only three time steps, which dominate the Causal Set path integral. They are therefore regarded as problematic, since such spacetimes do not resemble our extended universe. It remains an open question whether the continuum limit of the random matrix model (3.10) presented here retains these pathological configurations. We expect this analysis can be done using FRG techniques similarly to [19, 12].
Acknowledgements